The given function is f(x) = 2x^2 and g(x) = 3x+1
Find:
1.) f(x)^2
2.) f(x)^3-g(x)+g(x)^2

Answer by CubeyThePenguin(3113)

1) f(x)^2 = (2x^2)^2 = 4x^4

2) f(x)^3 - g(x) + g(x)^2 = (2x^2)^3 - (3x+1) + (3x+1)^2

= 8x^6 - 3x - 1 + 9x^2 + 6x + 1
= 8x^6 + 9x^2 + 3x